OpenStudy (anonymous):

A farmer has cows and ducks on a truck. These are the only creatures on the truck. Altogether, there are 15 heads and 44 legs on the truck. How many cows are on the truck? How many ducks?

OpenStudy (anonymous):

7 cows 8 ducks

OpenStudy (anonymous):

How do you know ?

OpenStudy (anonymous):

x is the cows y is the ducks x+y = 15 cows have 4 legs ducks have 2 4x+2y=44 multiply the first equation by 2 2x+2y =30 4x+2y=44 -2x=-14 divide each side by -2 x=7 plug x=7 into the first equation, x+y=15 7+y=15 y=8 so there are 7 cows and 8 ducks

OpenStudy (anonymous):

Thanks ! :)

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

Thanks for the medal

OpenStudy (anonymous):

Your welcomed.